# Embodied Co-Design for Rapidly Evolving Agents: Taxonomy, Frontiers, and Challenges

### v1

@@ IV-D2 Developmental Evolution
Brodbeck et al. [brodbeck2015morphological]
2015
Modular
Robot
Real-World
EA (GA) CPG EA (GA) Vector Locomotion Task Reward
@@ V Testbeds: Applying Embodied Co-Design
Embodied co-design has evolved rapidly due to the availability of high-fidelity simulated platforms and increasingly advanced real-world fabrication techniques. These resources provide standardized environments, reproducible tasks, and practical feedback loops that facilitate the automated design and validation of embodied agents [long2025survey]. This section surveys representative benchmarks, datasets, and applications of ECD across both simulated and real-world domains.
@@ V-A1 Rigid-Body Simulators
Widely used robot simulators, including MuJoCo [todorov2012mujoco], Bullet [panerati2021learning], Unity [juliani2018unity, Veenstra2023], Gazebo [koenig2004design], Webots [michel2004cyberbotics], SAPIEN [xiang2020sapien], DART [lee2018dart], Rostok [borisov2022reconfigurable], and V-Rep [rohmer2013v], provide native support for standard robot description formats such as XML, MJCF, URDF, and SDF. This compatibility enables the rapid modification of morphologies and establishes these simulators as the foundation for numerous works in ECD, as summarized in Tables IVI. The convergence of robotics, reinforcement learning, and photorealistic rendering has further increased the demand for scalable, GPU-accelerated simulation. Platforms such as NVIDIA Isaac Gym and its extensions [makoviychuk2021isaac, kulkarni2023aerial, gu2024humanoid] enable massively parallel environments that support the simultaneous training of large sets of morphological variants with controllers [cheng2024structural, wang2024evolution, chen2023evolving, ringel2025text2robot, jin2025evolutionary, chen2024pretraining, du2025efficient, schaff2022n], significantly accelerating ECD pipelines.
While much of the existing ECD literature relies on non-differentiable simulators that only provide forward dynamics, a growing number of differentiable platforms supporting inverse dynamics have emerged. These tools—Dojo [howell2022dojo], Brax [freeman2021brax], Neuralsim [heiden2021neuralsim], Nimble [werling2021fast] (the differentiable version of DART), DiffRedMax [xu2021accelerated], and SDRS [ye2024sdrs]—facilitate gradient-based co-design. Based on these efforts, frameworks such as Genesis [Genesis] and MuJoCo Playground [mujoco_playground_2025] aim to unify differentiable physics, GPU acceleration, and integrated rendering within a single API, enabling end-to-end pipelines for the co-design of complex embodied agents with neural network–based controllers, as shown in Table VII.
@@ V-A2 Soft-Body Simulators
For soft-bodied agents, early platforms such as SOFA [coevoet2017software, deimel2017automated], VoxCAD [cheney2013unshackling, cheney2014electro, cheney2018scalable], and VoxCraft [kriegman2021fractals, li2024reinforcement] remain influential baselines due to their detailed physical models of deformable materials. Lightweight two-dimensional frameworks, such as 2D-VSR-Sim [ferigo2025totipotent], further enable rapid algorithmic prototyping. Recently, EvolutionGym [bhatia2021evolution] introduces a comprehensive suite of tasks along with a diverse material palette for voxel-based soft robots, which has since been widely adopted in ECD research [katayama2023ga, dongleveraging2024, harada2024lamarckian, zhao2024morphological, wang2023curriculum, wang2023preco, xie2025accelerating, zhao2025multi, liu2023rapidly, saito2024effective, zhao2025cross, davis2023subtract, tanaka2022co, mertan2024no, mertan2025controller, nadizar2025enhancing, song2025laser, song2024morphvae, liu2025morphology]. Complementary to voxel-based systems, several platforms integrate differentiable soft-body physics, including DiffTaichi [hu2019difftaichi], ChainQueen [hu2019chainqueen], and DiffAqua [ma2021diffaqua]. SoftZoo [wang2023softzoo] extends this direction by providing a differentiable benchmark across various terrains (e.g., flat ground, snow, and water) and task objectives (e.g., speed, turning, and path following). In parallel, Sorotoki [caasenbrood2024sorotoki] provides an integrated MATLAB toolkit for co-designing soft robots.
@@ V-A3 Embodied Tasks Simulators
Notably, recent specialized simulators for embodied intelligence, such as iGibson [li2022igibson], AI2-THOR [kolve2017ai2], Habitat-Sim [szot2021habitat], and LIBERO [liu2023libero], provide realistic and interactive 3D environments that integrate rich visual, physical, and semantic information. These platforms support the training and evaluation of increasingly complex embodied agents, including those with advanced control paradigms (e.g., Vision-Language-Action models) [zitkovich2023rt, kim2024openvla, bjorck2025gr00t] and various morphologies (e.g., humanoids). Furthermore, they provide challenging benchmarks for long-horizon tasks, which makes them well-suited foundations for advancing ECD research in the future.
V-B Datasets
In addition to the aforementioned simulation platforms, several recent ECD datasets have been developed to support the systematic analysis of co-design across diverse embodied agents. RoboCrafter-QA [chen2025large] provides comparative data for evaluating whether LLMs can link high-level task descriptions to low-level morphological and material choices in soft robotics. RoboDesign1M [le2025robodesign1m] offers a large-scale dataset of one million robotic design samples across multiple domains, enabling tasks such as design–image generation, design-centric visual question answering, and design–image retrieval.
@@ V-C Real-World Applications
While simulation facilitates rapid exploration, physical validation remains the ultimate test of ECD. Recent advances in fabrication, multi-material 3D printing, and embedded sensing have made it feasible to realize co-designed morphologies. In this section, we summarize recent real-world applications.
@@ VII Conclusion
The interplay among morphology, control, and task environment is fundamental not only for biological organisms but also for the development of intelligent embodied agents. In this survey, we review recent advances in embodied co-design (ECD) and organize them into a hierarchical taxonomy. This taxonomy provides a structured framework that enables researchers to contextualize current progress and identify opportunities for cross-disciplinary collaboration. We also review notable applications in both simulated and physical environments. Although current results indicate that co-designed agents outperform fixed-body baselines, many open questions remain unanswered. We highlight these challenges and discuss promising directions for future research. We hope that this survey serves as both a foundational reference for newcomers and a comprehensive resource for experts seeking deeper insights into ECD.
@@
[Uncaptioned image] Yuxing Wang (Member, IEEE) received the B.Eng. degree in communication engineering from Southwest Minzu University (SMU) in 2020, and the M.Eng. degree in electronic and information engineering from Tsinghua University (THU) in 2023. He is currently pursuing a Ph.D. degree in control science and engineering at THU. Before that, he was a research assistant with THU and a research intern with Tencent AI Lab. His research interests include evolutionary learning and embodied intelligence, with a particular focus on evolutionary reinforcement learning for control optimization, and embodied co-design for modular robotic systems.

### v2

@@ IV-C3 Large-Model-Based Generation
IV-D Open-Ended Co-Design
Open-ended co-design aims to achieve continual improvement by allowing an agent’s morphology, controller, and environment to co-evolve. By removing fixed terminal objectives, these systems foster sustained innovation, as agents and environments mutually generate new challenges and adaptations, as shown in Fig. 3 (d). We categorize open-ended co-design into two primary subtopics: 1) Brain–Body–Environment Co-Evolution and 2) Developmental Evolution.
IV-D1 Brain-Body-Environment Co-Evolution
This approach treats the task environment as a co-optimized component alongside morphology and control. Foundational work in this direction is the Paired Open-Ended Trailblazer (POET) framework [wang2019poet, wang2020enhanced]. POET maintains a population of environment–controller pairs, simultaneously mutating environments and evolving agents. Policies that succeed in one environment may be transferred to other contexts. Stensby et al. [stensby2021co] extended POET to jointly evolve morphologies and controllers in tandem with their environments. Their results indicate that learned environmental progressions yield greater morphological diversity than handcrafted curricula and reduce the risk of premature convergence.
Recent work has introduced RL-driven co-evolution of the brain, body, and environment. Ao et al. proposed MECE [ao2023curriculum], which formulates a curriculum where both morphology and the environment evolve according to learned RL policies. This approach produces progressively challenging and structured tasks that promote sustained adaptation. Another emerging trend enhances environment evolution through the use of generative models. For instance, LLM-POET [aki2024llm] replaces manual environment mutation with a LLM that generates increasingly complex environments for a population of soft-robot agents. As agents improve, the LLM proposes novel tasks that escalate in difficulty, creating an open-ended loop of continual challenge and innovation.
IV-D2 Developmental Evolution
Developmental co-evolution draws inspiration from evolutionary developmental biology (Evo-Devo) and developmental robotics, where genotypes encode growth processes that shape both morphology and control over an agent’s lifetime [kriegman2018morphological, van2021comparing, arita2016alife, bongard2011morphological, bongard2010utility]. Early research has investigated developmental encodings, which include L-systems [rieffel2014growing], grammar rules, and neural networks [pontes2022single, joachimczak2016artificial]. Bongard [bongard2003evolving] demonstrated that artificial ontogeny can generate hierarchical, repeated structures and controllers that emerge as bodies grow, yielding more structured solutions.
The other research focus is on morphological development. Here, developmental rules adjust parameters, such as leg angles or actuator amplitudes, during learning [naya2020experiment] or progressively “unfreeze” degrees of freedom to facilitate skill acquisition [lungarella2002adaptivity, berthouze2004motor, savastano2012incremental]. Later work introduced mechanisms for growing new body parts during an agent’s lifetime [benureau2022morphological, jaafar2025eco, vujovic2017evolutionary], providing additional adaptive flexibility. For comprehensive overviews of developmental evolution, we refer readers to [naya2021morphological, wang2021survey]. Overall, open-ended co-design is not yet a unified methodology but rather a collection of emerging ideas that leverage ontogeny, plasticity, and staged learning, and it awaits further exploration.
@@ V Testbeds: Applying Embodied Co-Design
Embodied co-design has evolved rapidly due to the availability of high-fidelity simulation platforms and increasingly advanced real-world fabrication techniques. These resources provide standardized environments, reproducible tasks, and practical feedback loops that facilitate the automated design and validation of embodied agents. This section surveys representative benchmarks, datasets, and applications of ECD across both simulated and real-world domains.
@@ V-A1 Rigid-Body Simulators
As summarized in Table VII, widely used robot simulators, including MuJoCo [todorov2012mujoco], Bullet [panerati2021learning], Unity [juliani2018unity, Veenstra2023], Gazebo [koenig2004design], Webots [michel2004cyberbotics], SAPIEN [xiang2020sapien], DART [lee2018dart], Rostok [borisov2022reconfigurable], and V-Rep [rohmer2013v], provide native support for standard robot description formats such as XML, MJCF, URDF, and SDF. This compatibility facilitates rapid modifications to morphologies and establishes these simulators as foundational tools for numerous works in ECD. The convergence of robotics and AI has further increased the demand for scalable, GPU-accelerated simulations. Platforms such as NVIDIA Isaac Gym and its extensions [makoviychuk2021isaac, kulkarni2023aerial, gu2024humanoid] enable massively parallel environments that support simultaneous training of large sets of morphological variants with controllers [cheng2024structural, wang2024evolution, chen2023evolving].
While much of the existing ECD literature relies on non-differentiable simulators that only provide forward dynamics, a growing number of differentiable platforms supporting inverse dynamics have emerged. These tools—Dojo [howell2022dojo], Brax [freeman2021brax], Neuralsim [heiden2021neuralsim], Nimble [werling2021fast] (the differentiable version of DART), DiffRedMax [xu2021accelerated], and SDRS [ye2024sdrs]—facilitate physics-based co-design. Building on these efforts, frameworks such as Genesis [Genesis] and MuJoCo Playground [mujoco_playground_2025] aim to unify differentiable physics, GPU acceleration, and integrated rendering within a single API.
@@ V-A2 Soft-Body Simulators
For soft-bodied agents, early platforms such as SOFA [coevoet2017software, deimel2017automated], VoxCAD [cheney2013unshackling, cheney2014electro, cheney2018scalable], and VoxCraft [kriegman2021fractals, li2024reinforcement] remain influential due to their detailed physical models of deformable materials. Lightweight two-dimensional frameworks, such as 2D-VSR-Sim [ferigo2025totipotent], further enable rapid algorithmic prototyping. Recently, EvolutionGym [bhatia2021evolution] introduces a comprehensive suite of tasks along with a diverse material palette for voxel-based soft robots, which has since been widely adopted in ECD research [katayama2023ga, dongleveraging2024, harada2024lamarckian, zhao2024morphological, xie2025accelerating, zhao2025multi, liu2023rapidly, saito2024effective, zhao2025cross, davis2023subtract, tanaka2022co, mertan2024no, mertan2025controller, nadizar2025enhancing, song2025laser, song2024morphvae, liu2025morphology]. Complementary to voxel-based systems, several platforms integrate differentiable soft-body physics, including DiffTaichi [hu2019difftaichi], ChainQueen [hu2019chainqueen], and DiffAqua [ma2021diffaqua]. SoftZoo [wang2023softzoo] advances this direction by providing a benchmark across various terrains (e.g., flat ground, snow, and water) and task objectives (e.g., speed, turning, and path following). In parallel, Sorotoki [caasenbrood2024sorotoki] provides an integrated MATLAB toolkit.
@@ V-A3 Embodied Tasks Simulators
Notably, recent specialized simulators for embodied intelligence, such as iGibson [li2022igibson], AI2-THOR [kolve2017ai2], Habitat-Sim [szot2021habitat], and LIBERO [liu2023libero], provide realistic and interactive 3D environments that integrate rich visual, physical, and semantic information. These platforms support the training and evaluation of increasingly complex embodied agents, including those with advanced control paradigms (e.g., Vision-Language-Action models [zitkovich2023rt, kim2024openvla, bjorck2025gr00t]) and various morphologies (e.g., humanoids). Furthermore, they provide challenging benchmarks for long-horizon tasks involving navigation, manipulation, and multi-step planning, which makes them well-suited foundations for advancing ECD research in the future.
@@
V-B Datasets
In addition to the aforementioned simulation platforms, several ECD datasets have been developed to support the systematic analysis of agent design. RoboCrafter-QA [chen2025large] provides comparative data to evaluate whether LLMs can effectively link high-level task descriptions to low-level morphological and material choices in soft robotics. RoboDesign1M [le2025robodesign1m] offers a large-scale dataset of one million robotic design samples across multiple domains, enabling tasks such as design–image generation, design-centric visual question answering, and design–image retrieval.
Fig. 4: Co-designed embodied agents in both virtual and real environments. Example images numbered from (a) to (p) are adapted from [sims1994competition, hornby2001evolving, shim2003generating, Veenstra2023, ma2021diffaqua, ha2019reinforcement, luck2020data, zhao2020robogrammar, gupta2021embodied, xu2021accelerated, cheney2013unshackling, zardini2021seeking, medvet2021biodiversity, wang2023curriculum, wang2023softzoo, yue2025toward], respectively. Example images numbered from (q) to (y) are adapted from [nygaard2021environmental, revolve2, Strgar-RSS-24, xie2024map, schaff2022soft, liu_voxcraft_2020, kobayashi2024computational, matthews2023efficient, cheng2024structural], respectively.
@@ V-C Real-World Applications
While simulation facilitates rapid exploration, physical validation remains the ultimate test of ECD. Recent advances in fabrication, 3D printing, and embedded sensing have made it feasible to realize co-designed morphologies. In this section, we summarize recent real-world applications (Fig. 4).
@@ VII Conclusion
The interplay among morphology, control, and task environment is fundamental not only for biological organisms but also for the development of intelligent embodied agents. In this survey, we review recent advances in embodied co-design (ECD) and organize them into a hierarchical taxonomy. This taxonomy provides a structured framework that enables researchers to contextualize current progress and identify opportunities for cross-disciplinary collaboration. We also review notable applications in both simulated and physical environments. Although the current results indicate that co-designed agents outperform fixed-body baselines, many questions remain unanswered. We highlight these challenges and discuss promising directions for future research. We hope this survey serves as both a foundational reference for newcomers and a comprehensive resource for experts seeking deeper insights into ECD.
@@
[Uncaptioned image] Yuxing Wang (Member, IEEE) received the B.Eng. degree in communication engineering from Southwest Minzu University (SMU) in 2020, and the M.Eng. degree in electronic and information engineering from Tsinghua University (THU) in 2023. He is currently pursuing a Ph.D. degree in control science and engineering at THU. Before that, he worked as a research assistant at THU and as a research intern at Tencent AI Lab. His research interests include evolutionary learning and embodied intelligence, with a particular focus on evolutionary reinforcement learning for control optimization, and embodied co-design for modular robotic systems.
[Uncaptioned image] Zhiyu Chen received the B.Sc. degree in electronic engineering from Tsinghua University, Beijing, China, in 2025. He is currently pursuing a M.Eng. degree in artificial intelligence at Tsinghua University. His research interests are in reinforcement learning theory, embodied intelligence, and large language models.
[Uncaptioned image] Tiantian Zhang received the B.Sc. degree in automation from the Department of Information Science and Technology, Central South University, Changsha, China, in 2015, and the M.Sc. and Ph.D. degrees in control science and engineering from the Department of Automation, Tsinghua University, Beijing, China, in 2018 and 2024, respectively. She is currently a Post-Doctoral Researcher with the Center for Artificial Intelligence and Robotics, Shenzhen International Graduate School, Tsinghua University. Her research interests include data science, intelligent decision-making, and reinforcement learning. She has been serving as reviewers for Nature Communications, TNNLS, TMLR and AAAI.
[Uncaptioned image] Qiyue Yin received the Ph.D. degree from the National Laboratory of Pattern Recognition, Institute of Automation, Chinese Academy of Sciences (CASIA), Beijing, China, in 2017. He is currently an Associate Professor at CASIA. His major research interests include machine learning, pattern recognition and artificial intelligence on games.
[Uncaptioned image] Yongzhe Chang received the Ph.D. degree in machine learning&data analysis from University of New South Wales, Australia, in 2020. He was a post-doctoral research fellow with the AI&Robot laboratory, Tsinghua University, from 2021 to 2023. Then he joined Tsinghua University as a associate researcher since 2023. His research interests include machine learning, reinforcement learning, large model and intelligent robot, etc.
[Uncaptioned image] Zhiheng Li (Member, IEEE) received the Ph.D. degree in control science and engineering from Tsinghua University, Beijing, China, in 2009. He is currently an Associate Professor with the Department of Automation, Tsinghua University, and with the Graduate School at Shenzhen, Tsinghua University, Shenzhen, China. His research interests include traffic operation, advanced traffic management system, urban traffic planning, and intelligent transportation systems. Dr. Li is an Associate Editor for IEEE TRANSACTIONS ON INTELLIGENT TRANSPORTATION SYSTEMS.
[Uncaptioned image] Liang Wang (Fellow, IEEE) received the B.Eng. and M.Eng. degrees from Anhui University in 1997 and 2000, respectively, and the Ph.D. degree from the Institute of Automation, Chinese Academy of Sciences (CASIA) in 2004. From 2004 to 2010, he was a research assistant with Imperial College London, United Kingdom, and Monash University, Australia, a research fellow with the University of Melbourne, Australia, and a lecturer with the University of Bath, United Kingdom, respectively. He is currently a full professor of the Hundred Talents Program with the National Lab of Pattern Recognition, CASIA. He has widely authored or coauthored in highly ranked international journals, such as TPAMI and TIP, and leading international conferences, such as CVPR, ICCV, and ICDM. His research interests include machine learning, pattern recognition, and computer vision. He is an IAPR fellow.
[Uncaptioned image] Xueqian Wang (Member, IEEE) received the M.Sc. and Ph.D. degrees in control science and engineering from the Harbin Institute of Technology (HIT), Harbin, China, in 2005 and 2010, respectively. From June 2010 to February 2014, he was a Post-Doctoral Researcher with HIT. From March 2014 to November 2019, he was an Associate Professor with the Division of Informatics, Shenzhen International Graduate School, Tsinghua University, Shenzhen, China. He is currently a Professor and the Leader of the Center for Artificial Intelligence and Robotics, Shenzhen International Graduate School, Tsinghua University. His research interests include robot dynamics and control, teleoperation, intelligent decision-making and game playing, and fault diagnosis.